OpenStudy (anonymous):

]One solution of the equation 3x^2 − 16x + 5 = 0 is 1/3. What is the other solution?

OpenStudy (campbell_st):

well it factors to (3x - 1)(x - 5) = 0 just find the solutions..

OpenStudy (anonymous):

wait. wouldn't the 3 be outside of the parentheses, and doesn't the factors have to be added up to -16?

OpenStudy (campbell_st):

well just distribute and see if you get the original equation...

OpenStudy (anonymous):

*looks down at the ground* :3 maybe i got back to the original problem, maybe i didn't.. :3. How come math people are so smart??? @campbell_st

OpenStudy (anonymous):

so the other solution is x = -5?

OpenStudy (campbell_st):

its all about practice and identifiying... here is how I factored this equation works everytime if you have a quadratic \[ax^2 + bx + c = 0\] start with a*c in your question 3 times 5 = 15 find factors that add to -16, -15 and -1 then you set up the brackets for the binomials (ax + factor 1)(ax + factor 2) (3x -15)(3x -1) -------------------------- = -------------- =0 a 3 check for common factors 3(x - 5)(3x -1) -------------- = 0 3 a factor of 3 cancels... this will happen every time leaving (x -5)(3x -1) = 0 so you need to solve x - 5 = 0 3x - 1 = 0 to find the solutions to the equation.

OpenStudy (anonymous):

so the solutions are 1/3 and -5?

OpenStudy (campbell_st):

not -5 x - 5 = 0 x = 5 makes the equation true.

OpenStudy (anonymous):

oops meant +5*